# Studying stellar populations in Omega Centauri with phylogenetics

### v1

@@ 4.3 Connecting branches and populations through common stars
The interpretations regarding the nature of the populations found in the Optical sample depend on the dataset and the selection of stars, but using a different dataset does not contradict our results. We can reach that conclusion by studying the groups of stars in common. The trees in Fig. 8 show that purple stars in the Optical typically appear alongside yellow and red stars in the Infrared, while orange stars in the Optical are either yellow or blue infrared. Looking at the positions of these stars in Fig. 7 and 17 allows us to conclude that the orange-purple Optical branch, as well as the yellow-red-magenta branch in the Infrared, correspond to Block-Cen, where stars show signatures of in-situ star formation.
@@
GC-Cen stars, dominated mainly by orange stars in the Optical, are divided into blue and red stars in the Infrared. The GMM model prefers two groups for this population, which are divided in two branches in the Infrared tree. While the metallicity range is comparable in these branches, the O-Al anticorrelation is quite different. Indeed, among the three stars in the bottom red branch of the Infrared sample which are included in the Optical sample, SIR_61 is colored purple, while SIR_19 and SIR_45 are colored gray. These three stars are actually in the Child-Cen branch in the Optical tree, but SIR_45 is mixed with orange, and SIR_61 is at the bottom of the gray branch. We could thus attribute the red branch to Child-Cen instead of GC-Cen, but the metallicity trend and distribution are quite narrow. The Infrared Stellar Population Barplot shows that the blue, magenta and yellow groups are more related to each other than the red, suggesting that indeed the red is a more independent population, such as GCs falling into the potential well of the NSC.
@@
It might be possible that the Infrared sample has other Child-Cen. We need to consider that we do not have the same information for both trees. We have abundances of different species, which are measured using various methods and data. Furthermore, we have only 20% of stars in common between both samples. If the Infrared dataset includes Child-Cen, it does not cover a large range in s-process and metallicity to be assigned to a different ’caterpillar’ branch or a new GMM group. It would be worthwhile to see if some of the most metal-rich stars observed in Mészáros et al. (2021) indeed follow the chemical trends of the Optical sample. To do so, Optical spectroscopic follow-up observations are needed.
@@
It is worth commenting on the recent discussions about the anti-correlations of $\omega$ Cen, which tend to become weaker with higher metallicity (Alvarez Garay et al., 2024; Mészáros et al., 2021). The Child-Cen anti-correlation we observe for the highest metallicity stars could be the consequence of a channel that is different from the MgAl nucleosynthetic chain of globular clusters (Alvarez Garay et al., 2024). Indeed, Romano et al. (2023) studied various chemical evolution models to understand the chemical pattern of Terzan 5, which was found to host multiple populations similar to $\omega$ Cen, and is in the bulge of the Milky Way. In such a dense environment, the [Fe/H] spread, the age spread, and the chemical pattern of Terzan 5 is used to conclude that Terzan 5 can have multiple populations only because of in-situ star formation. Through a combination of star formation bursts with different duration and efficiency (which is dependent on the temperature of the gas, which might further depend on the temperature of the surroundings) and the ability to retain remnant stars for later Fe-peak production from SNe-Ia, it is possible to have an impact on the relative abundance ratios in such systems.
@@
Under the hypothesis that $\omega$ Cen is a NSC, we are able to identify both its main formation channels, namely GC accretion and in-situ formation (Neumayer et al., 2020). Using Fig. 6 of Fahrion et al. (2021), which relates the formation channels and the NSC progenitor galaxy’s mass, we conclude that the progenitor galaxy of $\omega$ Cen could have a mass of about $10^{9}\mathrm{M}_{\odot}$. This mass has been measured for the galaxy of the largest major merger event (Das et al., 2020), namely GES (Helmi et al., 2018). Our analysis thus provides evidence to support the claim of Pfeffer et al. (2021), that $\omega$ Cen might be the NSC of the GES.
@@ 4.4 The history of the formation and evolution of $\omega$ Cen
Our phylogenetic analysis using NJ trees allows us to study how our GMM groups are distributed and related to each other, enabling us to deduce that $\omega$ Cen appears to have formed via multiple formation channels: accretion of globular clusters, accretion of gas and stars belonging to primordial building blocks of galaxies, as well as in-situ star formation. GC-like stars belonging to G_O1-orange, are the most metal-poor and probably the most ancient ones in our dataset. Furthermore, the [La/Fe] abundances of these stars are low compared to the rest of the sample, and have a narrow distribution as well. This is consistent with a population formed rapidly and early on. They might be the relics of GC accretion into $\omega$ Cen. Because GC-Cen is mono-metallic, with a rather small dispersion, our data does not support a contribution of a large number of GCs contributing to the formation of this NSC. The other stellar populations in $\omega$ Cen show signatures of being the result of prolonged star formation. These populations belong mostly to G_O2-gray and G_O3-purple but some belong to G_O1-orange and are separated in two ‘caterpillar’ branches in the NJ tree.
@@
The metallicity distribution of Block-Cen, as well as its [O/Fe] and [La/Fe] trends, are consistent with chemical evolution models of Romano et al. (2023), which are specifically designed to replicate abundance patterns in dense and old environments, in their case, Terzan 5. Among the various abundance ratios, the [Al/Fe] enhancement has been found in potential ancient building blocks of the Milky Way (Malhan & Rix, 2024; Buckley et al., 2024). The fact that Block-Cen is enhanced in [Al/Fe] might thus be the result of a primordial population formed at the center of $\omega$ Cen’s progenitor galaxy which was accreted early on to form the NSC along with the accreted globular clusters. It is possible that both stars and gas were accreted into $\omega$ Cen.
@@
The second population, Child-Cen, corresponds to the branch with gray stars. Its [Fe/H] as well as [La/Fe] reach values that are very enhanced. [Al/Fe] is also very enhanced in this population, whereas the [O/Fe] values are very depleted. The low [O/Fe] might be a signature of gas accretion into $\omega$ Cen. While a combination of star formation bursts and interactions with the surrounding gas of a dense environment might yield an abundance pattern like the one found here, the pattern of a population like Child-Cen has not been intentionally modeled so far. The abundance trends of this branch are not like those observed in globular clusters, dwarf galaxies or in the populations of the Milky Way as discussed above, therefore we attribute it to a population that formed in-situ. Our phylogenetic analysis has allowed us to discover abundance trends that would require new models to explain them.
@@ 5 Concluding Remarks
Our inter-disciplinary methodology is able to provide more sophisticated analyses of the astrophysical evolution of a complex stellar system such as $\omega$ Cen than has been viable using traditional purely-astronomy approaches. Our interpretation that $\omega$ Cen might have built up from the merging of at least one metal-poor GC and more metal-rich populations whose broad metallicity distribution could be the result of prolonged star formation agrees well with the recent conclusions of Alvarez Garay et al. (2024) and the overall recent literature about the theoretical connection between $\omega$ Cen and NSCs (Brown et al., 2018; Pfeffer et al., 2021; Gray et al., 2024). Alvarez Garay et al. (2024) claim to reach their conclusions because of the improved precision in their abundance measurements compared to Johnson & Pilachowski (2010). Here, however, we use these seemingly less-precise datasets, but are still able to reach similar conclusions to Alvarez Garay et al. (2024). This is thanks to the novel methods we employ, adapted from evolutionary biology which has an established history in performing population analysis reconstructions
Nuclear star clusters might result from GC mergers, accretion of ancient building blocks, and in-situ star formation (Neumayer et al., 2020). Our analysis presents new evidence that $\omega$ Cen in favor of its nature being a nuclear star cluster. We find that its stellar populations have different chemical abundance pathways and relationships. The chemical pattern of the various branches in our NJ trees suggest that, in contrast to regular galaxies or genuine globular clusters, the $\omega$ Cen chemical history has been impacted by other processes, such as strong accretion and mixing. The multiple-channel formation scenario of NSCs can indeed account for the large variety of abundance patterns observed in $\omega$ Cen. Our phylogenetic analysis enables us to navigate the multidimensional chemical dataset produced by these formation pathways and disentangle the contributions of each channel. Furthermore, our analysis allows us to connect all these populations and reconstruct a possible shared history.
The presence of more than one ‘caterpillar’ branch in our tree thus rejects the hypothesis that $\omega$ Cen had an isolated evolutionary history. In fact, our branches could be associated with stellar populations of different origins: (1) GC-Cen: a globular cluster population; (2) Block-Cen: a stellar population that experienced star formation for more than 1 Gyr, with similar properties and perhaps origins to ancient clumpy populations found in dense environments such as the center of our Milky Way. (3) Child-Cen: a population with in-situ star formation, which could have formed inside $\omega$ Cen, but its chemical pattern has not been seen before so explaining it requires tailored modelling. Weather these are the true and only origins of the stellar populations of $\omega$ Cen, however, remains to be seen with more data, high resolution simulations of NSC, or with more detailed spectroscopic analysis of other NSC.
@@
Acknowledgements
This paper accomplishes one of the main scientific objectives of the interdisciplinary project Millennium Nucleus for the Evolutionary Reconstruction of the InterStellar medium (ERIS NCN2021_017). This work has been supported through student thesis fellowships of P.S, and K.W, in addition to funding for travel for P.J and K.W to visit or invite P.D, R.Y and X.H in several occasions. P.J. and C.A.G. thank Sven Buder, Joshua Povick and Nicole Buckley for fruitful suggestions and exploratory analyses on this paper. P.J. and F.C warmly thank Macarena Concha for introducing us to each other. P.J furthermore thanks FONDECYT REGULAR 1231057. C.A.G. acknowledges support from FONDECYT Iniciación 11230741. P.V and P.C acknowledge Milenio – ICN17_022 and NCN2024_040. P.D is supported by a UKRI Future Leaders Fellowship (grant reference MR/S032223/1). S.V. thanks ANID Beca Doctorado Nacional, 21220489). P.B.T acknowledges FONDECYT Regular 1240465. P.B.T and E.J. acknowledge support from ANID Basal Project FB210003. This project has received funding from the European Union Horizon 2020 Research and Innovation Programme under the Marie Sklodowska-Curie grant agreement No 734374-LACEGAL. A.R.A. acknowledges support from DICYT through grant 062319RA FONDECYT Regular 1230731.
@@ Appendix A Stars in common between Optical and Infrared sample
Table 1 lists the stars in common between both samples.
@@ B.1 PCA
Figure 9 illustrates the PCA outcomes, with the Optical sample displayed in blue on the left and the Infrared sample in red on the right. The histograms show the percentage of the variances of each principal component (dimension). We first focus on the Optical sample, which has abundance ratios of 11 elements. While the first 3 components explain 67% of the variance, with the second and the third dimensions contributing to a similar percentage (between 15 and 18%), a total of 6 dimensions are needed to explain 87% of the variance in the chemical space.
@@
The situation is slightly different regarding the Infrared sample, which has fewer abundances measured than the Optical sample. The first two components already explain more than 60% of the variance, and the first three components explain 80% of it. We comment that even though the Infrared sample is more limited in several elements, the sample also includes iron-peak, $\alpha-$capture and neutron-capture elements. However, it lacks the heavy r-process neutron-capture elements. The percentages cannot be directly compared to the Optical sample, since they are normalized to their own dataset.
@@
For the Optical sample, we obtain that Na, La, Si, Ca, Ti and Fe are the elements that most contribute to the first component. O, Ni, Ca and Al contribute most to the second component, and Fe, Eu, Al, and Sc contribute to the third component. We thus find that elements coming from different nucleosynthetic paths are combined in each principal component. The Infrared sample shows a similar situation. The most influential elements in the first component are O, Mg and Ce, in the second component are Al, Si and Ca, and the third component involves Fe, Ce and Ca.

### v2

@@ 4.3 Connecting branches and populations through common stars
The interpretations regarding the nature of the populations found in the Optical sample depend on the dataset and the selection of stars, but using a different dataset does not contradict our results. We can reach that conclusion by studying the groups of stars in common. The trees in Fig. 7 show that purple stars in the Optical typically appear alongside yellow and red stars in the Infrared, while orange stars in the Optical are either yellow or blue infrared. Looking at the positions of these stars in Fig. 6 and 16 allows us to conclude that the orange-purple Optical branch, as well as the yellow-red-magenta branch in the Infrared, correspond to Block-Cen, where stars show signatures of in-situ star formation.
@@
GC-Cen stars, dominated mainly by orange stars in the Optical, are divided into blue and red stars in the Infrared. The GMM model prefers two groups for this population, which are divided into two branches in the Infrared tree. While the metallicity range is comparable in these branches, the O-Al anticorrelation is quite different. Indeed, among the three stars in the bottom red branch of the Infrared sample which are included in the Optical sample, SIR_61 is colored purple, while SIR_19 and SIR_45 are colored gray. These three stars are actually in the Child-Cen branch in the Optical tree, but SIR_45 is mixed with orange, and SIR_61 is at the bottom of the gray branch. We could thus attribute the red branch to Child-Cen instead of GC-Cen, but the metallicity trend and distribution are quite narrow. The Infrared Stellar population bar plot shows that the blue, magenta and yellow groups are more related to each other than the red, suggesting that indeed the red is a more independent population, such as GCs falling into the potential well of the NSC.
@@
It might be possible that the Infrared sample has other stars belonging to Child-Cen. We need to consider that we do not have the same information for both trees. We have abundances of different species, which are measured using various methods and data. Furthermore, we have only 20% of stars that both samples have in common. If the Infrared dataset includes Child-Cen, it does not cover a large range in s-process and metallicity to be assigned to a different ”caterpillar” branch or a new GMM group. It would be worthwhile to see if some of the most metal-rich stars observed in Mészáros et al. (2021) indeed follow the chemical trends of the Optical sample. To do so, Optical spectroscopic follow-up observations are needed.
@@
It is worth commenting on the recent discussions about the anticorrelations of $\omega$ Cen, which tend to become weaker with higher metallicity (Alvarez Garay et al., 2024; Mészáros et al., 2021). The Child-Cen anticorrelation we observe for the highest metallicity stars could be the consequence of a channel that is different from the MgAl nucleosynthetic chain of globular clusters (Alvarez Garay et al., 2024). Indeed, Romano et al. (2023) studied various chemical evolution models to understand the chemical pattern of Terzan 5, which was found to host multiple populations similar to $\omega$ Cen, and is in the bulge of the Milky Way. In such a dense environment, the [Fe/H] spread, the age spread, and the chemical pattern of Terzan 5 is used to conclude that Terzan 5 can have multiple populations only because of in-situ star formation. Through a combination of star formation bursts with different duration and efficiency (which is dependent on the temperature of the gas, which might further depend on the temperature of the surroundings) and the ability to retain remnant stars for later Fe-peak production from SNe-Ia, it is possible to have an impact on the relative abundance ratios in such systems.
@@
Under the hypothesis that $\omega$ Cen is a NSC, we are able to identify both its main formation channels, namely GC accretion and in-situ formation (Neumayer et al., 2020). Using Fig. 6 of Fahrion et al. (2021), which relates the formation channels and the NSC progenitor galaxy’s mass, we conclude that the progenitor galaxy of $\omega$ Cen could have a mass of about $10^{9}\mathrm{M}_{\odot}$. This mass has been measured for the galaxy of the largest major merger event (Das et al., 2020), namely Gaia-Enceladus Sausage (GES Helmi et al., 2018; Belokurov et al., 2018). Our analysis thus provides evidence to support previous claims of Bekki & Freeman (2003) and Pfeffer et al. (2021), that $\omega$ Cen might be the NSC of the GES. Indeed, among all clusters that are dynamically associated with the GES, according to Massari et al. (2019), $\omega$ Cen is the one with the highest binding energy (see also the extensive discussion from Limberg et al., 2022, about this scenario).
We note that the connection between the GES and $\omega$ Cen does not need to imply a chemical similarity between the $\omega$ Cen populations and the field accreted stars that have been associated with GES by various works in the literature. Indeed, these stars are more $\alpha-$poor, Al-poor, and Eu-rich compared to $\omega$ Cen (see also Fig. 14 to see some elements Matsuno et al., 2021; Carrillo et al., 2022; Nissen & Schuster, 2011; Das et al., 2020; da Silva & Smiljanic, 2023). The chemical similarity between GES and $\omega$ Cen might be expected if $\omega$ Cen were a typical GC, such as NGC1261, which Massari et al. (2019) have attributed be part of GES. Indeed, Koch-Hansen et al. (2021) provided evidence of its chemical similarity with GES. But expecting stars from a NSC and the field of the parent galaxy sharing the same chemical composition might imply they share the same star formation history, which does not need to be true (Neumayer et al., 2020).
Let us furthermore consider the current belief that M54 is the NSC of the Sagittarius (Sgr) galaxy (Alfaro-Cuello et al., 2019). The chemical abundance analysis of M54 and the Sgr performed by Carretta et al. (2010) showed a continuation of $\alpha$ elements between M54 and the Sgr core, but the metallicity of both populations did not overlap. Abundances such as O, Na, Al differ due to the anticorrelations present in M54, which are not in the field. In Vitali et al. (2024), more metal-poor stars were included in the analysis of the field stars, overlapping better with the M54 metallicity coverage of Carretta et al. (2010). The overlap shows that $\alpha-$ abundances of Sgr’s core at a $\mathrm{[Fe/H]}\sim-1.5$ is of the order of 0.2 dex, while the abundances of M54 is of the order of $0.4-0.5$ dex, although this comparison is subject to uncertainties due to the different methods employed in these independent works. This difference is consistent with what is found between GES and $\omega$ Cen. More discussions about the chemical connection between GES stars and $\omega$ Cen and the role of europium can be further found in App. D.
@@ 4.4 The history of the formation and evolution of $\omega$ Cen
Our phylogenetic analysis using NJ trees allows us to study how our GMM groups are distributed and related to each other, enabling us to deduce that $\omega$ Cen appears to have formed via multiple formation channels: accretion of globular clusters, accretion of gas and stars belonging to primordial building blocks of galaxies, as well as in-situ star formation. GC-like stars belonging to G_O1-orange, are the most metal-poor and probably the most ancient ones in our dataset. Furthermore, the [La/Fe] abundances of these stars are low compared to the rest of the sample, and have a narrow distribution as well. This is consistent with a population formed rapidly and early on. They might be the relics of GC accretion onto $\omega$ Cen. Because GC-Cen is mono-metallic, with a rather small dispersion, our data does not support a contribution of a large number of GCs contributing to the formation of this NSC. The other stellar populations in $\omega$ Cen show signatures of being the result of prolonged star formation. These populations belong mostly to G_O2-gray and G_O3-purple but some belong to G_O1-orange and are separated in two ‘caterpillar’ branches in the NJ tree.
@@
The metallicity distribution of Block-Cen, as well as its [O/Fe] and [La/Fe] trends, are consistent with chemical evolution models of Romano et al. (2023), which are specifically designed to replicate abundance patterns in dense and old environments, in their case, Terzan 5. Among the various abundance ratios, the [Al/Fe] enhancement has been found in potential ancient building blocks of the Milky Way (Malhan & Rix, 2024; Buckley et al., 2024). The fact that Block-Cen is enhanced in [Al/Fe] might thus be the result of a primordial population formed at the center of $\omega$ Cen’s progenitor galaxy, which was accreted early on to form the NSC along with the accreted globular clusters. It is possible that both stars and gas were accreted onto $\omega$ Cen.
@@
The second population, Child-Cen, corresponds to the branch with gray stars. Its [Fe/H] as well as [La/Fe] reach values that are very enhanced. [Al/Fe] is also very enhanced in this population, whereas the [O/Fe] values are very depleted. The low [O/Fe] might be a signature of pristine gas accretion onto $\omega$ Cen (D’Ercole et al., 2012). While a combination of star formation bursts and interactions with the surrounding gas of a dense environment might yield an abundance pattern like the one found here, the pattern of a population like Child-Cen has not been intentionally modeled so far. The abundance trends of this branch are not like those observed in globular clusters, dwarf galaxies or in the populations of the Milky Way as discussed above, therefore we attribute it to a population that formed in situ. Our phylogenetic analysis has allowed us to discover abundance trends that would require new models to explain them.
As a final note, we point to the recent independent analysis of Mason et al. (2025), which was announced at the same time as this work. That work used a different dataset and different methods to reach a remarkably consistent conclusions about the three populations of $\omega$ Cen: P1 (our Block-Cen), P2 (our Child-Cen), and IM (our GC-Cen). While the similarities and differences of the exact properties and abundance patterns of these populations need to be further studied, it seems that new data and methods are converging into disentangling the populations of $\omega$ Cen and revealing its nature as a NSC.
@@ 5 Concluding remarks
Our interdisciplinary methodology is able to provide more sophisticated analyses of the astrophysical evolution of a complex stellar system such as $\omega$ Cen than has been viable using traditional, purely astronomy approaches. Our interpretation that $\omega$ Cen might have built up from the merging of at least one metal-poor GC and more metal-rich populations whose broad metallicity distribution could be the result of prolonged star formation agrees well with the recent conclusions of Alvarez Garay et al. (2024) and the overall recent literature about the theoretical connection between $\omega$ Cen and NSCs (Brown et al., 2018; Pfeffer et al., 2021; Gray et al., 2025). Alvarez Garay et al. (2024) claim to reach their conclusions because of the improved precision in their abundance measurements compared to Johnson & Pilachowski (2010). Here, however, we use these seemingly less precise datasets, but are still able to reach similar conclusions to Alvarez Garay et al. (2024). This is thanks to the novel methods we employ, adapted from evolutionary biology, which has an established history in performing population analysis reconstructions
[…]
The presence of more than one ‘caterpillar’ branch in our tree thus rejects the hypothesis that $\omega$ Cen had an isolated evolutionary history. In fact, our branches could be associated with stellar populations of different origins: (1) GC-Cen: a globular cluster population; (2) Block-Cen: a stellar population that experienced star formation for more than one Gyr, with similar properties and perhaps origins to ancient clumpy populations found in dense environments such as the center of our Milky Way. (3) Child-Cen: a population with in-situ star formation, which could have formed inside $\omega$ Cen, but its chemical pattern has not been seen before so explaining it requires tailored modeling. Whether these are the true and only origins of the stellar populations of $\omega$ Cen, however, remains to be seen with more data, high resolution simulations of NSCs, or with more detailed spectroscopic analysis of other NSCs.
@@
6 Data availability
Tables with abundance determination as introduced in Sect. 2 are only available in electronic form at the CDS via anonymous ftp to cdsarc.u-strasbg.fr (130.79.128.5) or via http://cdsweb.u-strasbg.fr/cgi-bin/qcat?J/A+A/.
Acknowledgements.
This paper accomplishes one of the main scientific objectives of the interdisciplinary project Millennium Nucleus for the Evolutionary Reconstruction of the InterStellar medium (ERIS NCN2021_017). This work has been supported through student thesis fellowships of P.S, and K.W, in addition to funding for travel for P.J and K.W to visit or invite P.D, R.Y and X.H in several occasions. P.J. and C.A.G. thank Keaghan Yaxley, Sven Buder, Joshua Povick and Nicole Buckley for fruitful suggestions and exploratory analyses on this paper. P.J. and F.C warmly thank Macarena Concha for introducing us to each other. P.J furthermore thanks FONDECYT REGULAR 1231057. C.A.G. acknowledges support from FONDECYT Iniciación 11230741. P.V and P.C acknowledge Milenio – ICN17_022 and NCN2024_040. P.D is supported by a UKRI Future Leaders Fellowship (grant reference MR/S032223/1). S.V. thanks ANID Beca Doctorado Nacional, 21220489). P.B.T acknowledges FONDECYT Regular 1240465. P.B.T and E.J. acknowledge support from ANID Basal Project FB210003. This project has received funding from the European Union Horizon 2020 Research and Innovation Programme under the Marie Sklodowska-Curie grant agreement No 734374-LACEGAL. A.R.A. acknowledges support from DICYT through grant 062319RA FONDECYT Regular 1230731. GG acknowledges support from the Leverhulme Trust under grant EM-2025-007. The authors warmly thank the referee for the constructive and respectful feedback.
@@ Appendix A Stars in common between Optical and Infrared sample
Table 1 lists the stars in common between both samples. The identification number from Gaia DR3, the APOGEE ID and the Leid ID are indicated for these stars, in addition to our own designation. We add the Gaia G apparent magnitude for reference.
@@ B.1 PCA
Figure 8 illustrates the PCA outcomes, with the Optical sample displayed in blue on the left and the Infrared sample in red on the right. The histograms show the percentage of the variances of each principal component (dimension). We first focus on the Optical sample, which has abundance ratios of 11 elements. While the first three components explain 67% of the variance, with the second and the third dimensions contributing to a similar percentage (between 15 and 18%), a total of six dimensions are needed to explain 87% of the variance in the chemical space.
@@
The situation is slightly different regarding the Infrared sample, which has fewer abundances measured than the Optical sample. The first two components already explain more than 60% of the variance, and the first three components explain 80% of it. We comment that even though the Infrared sample is more limited in several elements, the sample also includes iron-peak, $\alpha-$capture, and neutron-capture elements. However, it lacks the heavy r-process neutron-capture elements. The percentages cannot be directly compared to the Optical sample, since they are normalized to their own dataset.
@@
For the Optical sample, we obtain that Na, La, Si, Ca, Ti, and Fe are the elements that most contribute to the first component. O, Ni, Ca, and Al contribute most to the second component, and Fe, Eu, Al, and Sc contribute to the third component. We thus find that elements coming from different nucleosynthetic paths are combined in each principal component. The Infrared sample shows a similar situation. The most influential elements in the first component are O, Mg, and Ce, in the second component are Al, Si, and Ca, and the third component involves Fe, Ce, and Ca.